OpenStudy (anonymous):

f(x) = 2/3x - 6 and g(x) = 3x + 18/2 I got down to this answer 2/3(3x+18/2)-6 The final answer is "x" but i'm not sure how to get down to that answer. Help? :)

OpenStudy (mertsj):

\[\frac{2}{3}(\frac{3x+18}{2})-6=x+6-6=x\]

OpenStudy (phi):

this is easier to do if you put in parentheses: